OpenStudy (anonymous):

for which numbers \( \alpha \) is it true that \( x^2 + \alpha xy + y^2 > 0 \) whenever \( x \) and \( y \) are not both 0?

OpenStudy (jamesj):

Seriously, you should try and figure it out. Or are you truly completely stuck?

OpenStudy (anonymous):

I think we need to use the \( ax^2 + bx + c \) equation

OpenStudy (jamesj):

you can. or you can use all of the practice you just gained on completing the square.

OpenStudy (jamesj):

what happens for instance when alpha = 2?

OpenStudy (anonymous):

when alpha = 2, \[(x + y)^2 is >= 0\]

OpenStudy (jamesj):

try some experiments. Whan happens when alpha = 2, 1 or 0?

OpenStudy (jamesj):

so obviously alpha = 2 doesn't work.

OpenStudy (anonymous):

if alpha is 0 the expression is also > 0

OpenStudy (anonymous):

if alpha is -2 then the expression is > 0

OpenStudy (jamesj):

is it? I don't think so.

OpenStudy (anonymous):

when alpha is -2 then it becomes like \( x^2 - 2xy + y^2 \) which becomes \( (x-y)^2 \)which is positive (when not both x and y are 0)

OpenStudy (jamesj):

but is zero whenever x = y, for example x = y = 1.

OpenStudy (anonymous):

right.

OpenStudy (anonymous):

so it's not true at alpha = -2

OpenStudy (jamesj):

or alpha = 2.

OpenStudy (anonymous):

but it's true at 0

OpenStudy (jamesj):

try completing the square and see what happens. You know what you want: the sum of positive multiples of perfect squares

OpenStudy (anonymous):

\( x^2 + xy + y^2 + xy - xy = (x + y)^2 - xy\)

OpenStudy (anonymous):

this is when alpha is 1 if y = -x, then the thing in parentheses is 0 while the remaining term is positive since -x * -x = x^2

OpenStudy (anonymous):

I think same thing with alpha = -1

OpenStudy (jamesj):

When alpha = 1, write a for alpha x^2 + axy + y^2 = (x + y/2)^2 + (3/4).y^2

OpenStudy (anonymous):

alpha = 0?

OpenStudy (jamesj):

agd: try completing the square for the general case

OpenStudy (anonymous):

when alpha = b ?

OpenStudy (jamesj):

No. If I said to you: complete the square of this expression, x^2 + axy what would you do?

OpenStudy (anonymous):

Add y^2/4 and subtract y^2/4

OpenStudy (anonymous):

yeah I will do what GT said

OpenStudy (jamesj):

a^2y^2/4 yes, to obtain what expression?

OpenStudy (anonymous):

Actually a^2 to it.

OpenStudy (anonymous):

a positive expression :-D

OpenStudy (anonymous):

oh! the discriminant!

OpenStudy (anonymous):

(x+ay/2)^2 - a^2y^2/4

OpenStudy (jamesj):

wrong. It's not positive in general. x^2 + axy = (x+ay/2)^2 - a^2y^2/4

OpenStudy (jamesj):

right GT. Now, complete the square of the expression x^2 + axy + y^2 for general a.

OpenStudy (anonymous):

\[(x^2+a^2y^2/2)^2 + y^2 - a^2y^2/2 = y^2(1 - a^2/2) \]

OpenStudy (anonymous):

now it's clearer

OpenStudy (jamesj):

what you just wrote down isn't right, you know that, yes?

OpenStudy (anonymous):

yeah I wrote something wrong

OpenStudy (anonymous):

(x+y)^2+xy(a-2)

OpenStudy (jamesj):

\[ x^2 + axy + y^2 = (x+ay/2)^2 + (1-a^2/4)y^2 \] Now when is that positive?

OpenStudy (anonymous):

that's positive if and only if 1- a^2/4 is positive

OpenStudy (anonymous):

and y is not 0

OpenStudy (jamesj):

yes. We adopt everywhere to hypothesis that x and y are not both zero. And for what values of a is that true?

OpenStudy (anonymous):

-2 < a < 2

OpenStudy (anonymous):

thanks!

OpenStudy (jamesj):

yes

OpenStudy (anonymous):

It takes me half an hour to do these questions :(

OpenStudy (jamesj):

that's why it's called learning. This isn't 8th grade any more.

OpenStudy (anonymous):

right :-D

OpenStudy (anonymous):

I'll try the next one on my own and just post my work so folks can check if I messed something up.

OpenStudy (jamesj):

Not studying for the AI final?

OpenStudy (anonymous):

I still didn't watch the Natural Language Processing videos yet :(